2 For this problem, assume earned premium is the same as the written premium. Here is a part of financials for State Auto Financial
dollars in millions
2018 2017 2016 2015 2014
Earned Premiums     1,238.00     1,276.10     1,291.10     1,275.90     1,068.70
Loss Ratio 64.30% 72% 73% 67.80% 72%
Expense Ratio 36.30% 35.10% 33.50% 33.80% 34.50%
Dividend Ratio 0.0% 0.0% 0.0% 0.0% 0.0%
Combined Ratio
Loss and Loss Adjustment Expenses
Underwriting Expenses
Answer #1

Ans: For loss and adjustment expenses

Loss ratio* Earned premium

For 2018= 1,238*64.30%

= 796.03

For 2017= 1,276.10*72%

=918.79

For 2016= 1,291.10*73%

=942.50

For 2015= 1,275.90*67.80%

=865.06

For 2014= 1,068.70*72%

=769.46

For Underwriting Expense = Expense ratio* Earned premiums

For 2018= 1,238*36.30% = 449.39

For 2017= 1,276.10*35.10% = 447.91

For 2016= 1,291.10*33.50% = 432.51

For 2015= 1,275.90*33.80% = 431.25

For 2014= 1,068.70*34.50% = 368.70

For Combined ratio = Loss and adjustment expense+Underwriting expense)/Earned premiums

For 2018= (449.39+796.03)/1,238

= 100.59%

For 2017= (918.7+447.90)/1,276.10

= 107.09%

For 2016= (942.50+432.51)/1,291.10

= 106.50%

For 2015= (431.25+865.06)/1,275.90

= 101.59%

For 2014= (368.70+769.46)/1,068.70

= 106.49%